### v1

@@ Abstract
Utilizing the framework of free probability, we analyze the spectral and operator statistics of the Rosenzweig–Porter random matrix ensembles, which exhibit a rich phase structure encompassing ergodic, fractal, and localized regimes. Leveraging subordination formulae, we develop a perturbative scheme that yields semi-analytic expressions for the density of states up to second order in system size, in good agreement with numerical results. We compute higher-point correlation functions in the ergodic regime using both numerical and suitable analytic approximations. Our analysis of operator statistics for various spin operators across these regimes reveals close agreement with free probability predictions in the ergodic phase, in contrast to persistent deviations observed in the fractal and localized phases, even at late times. Notably, the fractal phase exhibits partial features of asymptotic freeness while retaining memory of the initial spectrum, highlighting the importance of non-localized eigenstates for the emergence of free probability behavior. Employing distance measures and statistical tools such as the $\chi^{2}$ statistic, Kullback–Leibler divergence, and Kolmogorov–Smirnov hypothesis testing, we define a characteristic time scale – the free time – that marks the onset of the validity of free probability predictions for operator spectral statistics in the ergodic phase. Remarkably, our findings demonstrate consistency across these different approaches.
@@ 1 Introduction
Although similar techniques have been previously applied to analyze the DOS in various models under constrained settings, our method has the practical advantage that it directly uses the free probability results (which otherwise one usually needs to establish first, e.g. using the replica technique and so on) and is also valid for a general class of Hamiltonians where one of the two terms of the Hamiltonian is Gaussian random matrix. As one of the main results of the paper, we obtain the first two orders corrections to the Gaussian distribution of the DOS of the RP model when the system is in the fractal phase. By numerical analysis, we show that our approximate analytical expression for the DOS is in very good agreement with it.
We further study the statistics of the sum of two deterministic spin operators, one of which is time-evolved by the RP model Hamiltonian. At the initial stage, the eigenvalue distribution of the total operator adheres to the free binomial distribution. However, as time progresses, the eigenvalue distribution exhibits significant deviations from this initial distribution. Specifically, in the ergodic regime, the eigenvalue statistics of the operator sum show a transition from the multinomial distribution to the one predicted by the free probability distribution when the component operators are free. The multinomial distribution (discussed later) is the general case of the binomial distribution, which holds only for spin-half operators. Conversely, in the fractal and localized regimes, the eigenvalue distributions diverge substantially from the predictions obtained from free probability theory. This observation lends credence to the broader assertion that chaos leads to the asymptotic freeness between two operators at late times.
@@ 2 Elements of Free Probability theory
Having discussed the condition that two mutually free random variables, say $a$ and $b$, should satisfy [^3] Throughout this paper, we assume that these two variables are self-adjoint and have well-defined eigenvalue density., we now describe the procedure for obtaining the free probability prediction for the density of eigenvalues of the sum ($a+b$) of these two operators, which will be used in Sec. 6 to compare with numerically obtained distributions. The eigenvalue distribution of $a+b$ is called the free convolution of the eigenvalue distributions $\rho_{a}(\lambda)$ and $\rho_{b}(\lambda)$, and is usually denoted as $\rho_{a+b}(\lambda)=\rho_{a}(\lambda)\boxplus\rho_{b}(\lambda)$ nica2006lectures .
The first ingredient we need to define is the resolvent operator associated with the operators $a$ and $b$, which is the Cauchy transformation of the DOS of the respective operator. E.g.,
@@ 2.1 Summary of Known Results on Free Additive Convolution
For operators with spin larger than unity, it becomes increasingly challenging to obtain analytic formulae for the free additive convolution of the sum of operators, due to the fact that one needs to solve order-$s$ algebraic equations. Fortunately, one can obtain such distributions numerically using subordination function methods; see, for instance speicher2019lecturenotesfreeprobability . Such methods were applied in Camargo:2025zxr for the derivation of the free additive convolution of spin-3/2 and spin-2 operators. Figure 1 shows the respective distribution for spin $1/2$, spin $1$, and spin $3/2$. We review the subordination function method in Appendix B.
@@ 2.2 Asymptotic Freeness Produced by Chaotic Dynamics
Given two operators $A$ and $B$ that are not initially free, the time evolution of $B$ under a chaotic Hamiltonian is expected to produce a Heisenberg operator $B(t)=e^{iHt}Be^{-iHt}$ that becomes free from $A$ Jindal:2024zcg ; Chen:2024zfj . This emergence of freeness defines a novel notion of quantum chaos, which can be incorporated into the quantum ergodic hierarchy Camargo:2025zxr .
Freeness arises only in infinite-dimensional Hilbert spaces. When operators $A$ and $B$ are represented as $N\times N$ matrices, freeness can emerge only in the limit $N\to\infty$, a regime associated with asymptotic freeness. For finite $N$, cumulants never vanish completely and retain residual values at late times. However, under chaotic dynamics – i.e., when the Hamiltonian exhibits features akin to random matrix theory – the magnitude of these residual cumulants decreases with increasing $N$, suggesting their vanishing in the $N\to\infty$ limit. In this sense, the finite-$N$ system “knows” about its large-$N$ counterpart.
This connection also manifests in operator statistics. Although true freeness between $A$ and the time-evolved operator $B(t)$ requires infinite $N$, the spectral statistics of $A+B(t)$ at finite $N$ often agree with free probability predictions. This agreement serves as a practical signature – or “smoking gun” – of asymptotic freeness.
@@ 4.1 Cauchy and R-transforms
From the functional inverse of the resolvent of an operator $A$ (i.e., the Cauchy transformation of the density of eigenvalues of the operator $A$), we define the $R$-transformation
$G_{A}^{-1}(z)\equiv\mathcal{B}_{A}(z)=\frac{1}{z}+R_{A}(z)~{}.$ (20)
We note that the expression for the resolvent of the operator $A$ in terms of its DOS, $\rho_{A}(\lambda)$ as
$G_{A}(z)=\int\frac{\rho_{A}(\lambda)}{z-\lambda}~{}\text{d}\lambda~{}.$ (21)
Here, the integral is over the support of the distribution $\rho_{A}(\lambda)$. Note that $G_{A}(z)$ is essentially the Cauchy transformation of $\rho_{A}(\lambda)$.
When the operator $A$ is the Hamiltonian of some quantum mechanical system belonging to some random matrix ensemble, one can relate the resolvent with the time evolution operator as follows. Consider the trace of the ensemble-averaged time evolution operator,
@@
$\rho_{H}(\lambda)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\text{d}\lambda~{}e^{-itH}~{}\mathcal{U}_{H}(t)~{}.$ (23)
@@ 4.2 Addition of two free operators: subordination formulae for the resolvent
Now, for addition of two such operators $A$ and $B$, which are free, from the additivity of the $R$ transformation, we have, for the operator $C=A+B$, the following,
@@
$\mathcal{B}_{A}\big{(}G_{C}(z)\big{)}=z-R_{B}\big{(}G_{C}(z)\big{)}\,.$ (26)
Finally, applying $G_{A}$ on both sides, we get a self-consistency equation for $G_{C}(z)$,
$G_{C}(z)=G_{A}\big{(}z-R_{B}\big{(}G_{C}(z)\big{)}\big{)}\,.$ (27)
One can perform an entirely analogous procedure for the operator $B$ and obtain
$G_{C}(z)=G_{B}\big{(}z-R_{A}\big{(}G_{C}(z)\big{)}\big{)}~{}.$ (28)
In the following, we use these formulae to obtain an expression for the DOS for the RP model. The choice between the two formulae for obtaining the density of eigenvalues of $C$ – which can be extracted from $G_{C}(z)$ using the Stieltjes inversion formula – depends on whether the resolvent and the $R$-transformation are more straightforward for the operators $A$ or $B$.
Before proceeding further, we note that a self-consistency equation analogous to the one in (27) for the resolvent of the RP model was obtained in Venturelli:2022hka using the replica trick. The authors also discussed the connection of their formula with the one obtained from the law of addition in free probability. The discussion on the approximate evaluation of the DOS of the RP model below follows closely the discussion in Venturelli:2022hka . However, we have kept the discussion below general enough so that it applies not only for the RP model, but addition of two non-commutative operators, one of which has the Wigner semi-circle as the DOS.
@@ 4.3 A perturbative scheme for the resolvent
For most of the choices of the operators $A$ and $B$, the expressions for the corresponding resolvents are sufficiently complicated so that it is difficult to solve either of the equations in eqs. (27) or (28) for $G_{C}(z)$. Therefore, we now proceed as follows to obtain a perturbative expression for the DOS of an operator of the form $C=A+\alpha B$, where $A$ and $B$ are mutually free and the expressions for their DOS are assumed to be known. Here $\alpha$ is a parameter which we assume to be small so that one can set up a perturbation series with respect to this parameter. E.g., for the RP model Hamiltonian, $\alpha=N^{-\gamma/2}$. Using the scaling transformation rule of the $R$-transformation [^7] The $R$-transformation of the density of eigenvalues of an operator $A$ follows the scaling law: $R_{\alpha A}(z)=\alpha R_{A}(\alpha z)$, where $\alpha$ is a constant., and the definition (in eq. (21)), we have, from eq. (27),
$G_{C}(z)$ $=\int\text{d}\lambda~{}\frac{\rho_{A}(\lambda)}{z-\alpha R_{B}\big{(}\alpha G_{C}(z)\big{)}-\lambda}$
$=i\int_{0}^{\infty}~{}\text{d}u~{}\chi_{A}(-u)\exp[-iu\Big{(}z-\alpha R_{B}\big{(}\alpha G_{C}(z)\big{)}\Big{)}\Big{]}~{}.$ (29)
Here we have defined the characteristics function $\psi_{A}(u)$ of the distribution $\rho_{A}(\lambda)$ by a Fourier transformation
$\chi_{A}(u)=\int\text{d}\lambda~{}\rho_{A}(\lambda)~{}e^{-i\lambda u}~{}.$ (30)
Comparing this expression with the inverse of the equation. (23), we see that the characteristic function is nothing but the ensemble-averaged evolution operator generated by $A$, i.e., $\chi_{A}(-t)=\mathcal{U}_{A}(t)$.
@@
$G_{C}(z)=i\int_{0}^{\infty}~{}\text{d}u~{}\chi_{A}(-u)~{}e^{-iuz}\sum_{n}\frac{(iu\alpha)^{n}}{n!}\big{(}R_{B}\big{(}\alpha G_{C}(z)\big{)}\big{)}^{n}~{}.$ (31)
Writing in terms of the free cumulants ($\kappa_{j}(B)$) of the operator $B$, this has the form
$G_{C}(z)=\sum_{n}\frac{i^{n+1}\alpha^{n}}{n!}\Big{(}\sum_{j=1}^{\infty}\kappa_{j}(B)\big{(}\alpha G_{C}(z)\big{)}^{j}\Big{)}^{n}\int_{0}^{\infty}~{}\text{d}u~{}\chi_{A}(-u)~{}u^{n}~{}e^{-iuz}~{}.$ (32)
So far, this expression for $G_{c}(z)$ is exact. For a general operator $B$, its $R$-transformation can be an arbitrary function of $z$, so that it is difficult to obtain a perturbation series with $\alpha$ as a small parameter. However, one particular case where the above expression becomes easily tractable is when the operator $B$ has Wigner semi-circle as the DOS (which is the case also for the RP model considered in this paper), i.e., $\rho_{B}(\lambda)=\frac{1}{2\pi}\sqrt{4-\lambda^{2}}$. In that case [^8] The Wigner semi-circle is the analogue of the Gaussian distribution for non-commutative probability, i.e., this distribution has only the first free cumulant non-zero. Since $R(z)$ is the generator of the free cumulants, this is equal to $R(z)=z$ for the semi-circle distribution. $R(z)=z$, and the expression in eq. (31) (or (32)) simplifies to
$G_{C}(z)$ $=\sum_{n}\frac{i^{n+1}}{n!}\alpha^{2n}\big{(}G_{C}(z)\big{)}^{n}\int_{0}^{\infty}\text{d}uu^{n}\chi_{A}(-u)e^{-iuz}\,,$
$=\sum_{n}\frac{(-1)^{n}}{n!}\alpha^{2n}\big{(}G_{C}(z)\big{)}^{n}\partial_{z}^{n}G_{A}(z)\,,$ (33)
where to obtain the last expression, we have used an analogous relation to that of the one in (24), valid for the operator $A$.
Evaluating the terms of the above sum recursively, one can obtain the resolvent $G_{C}(z)$ with corrections of different orders of $\alpha$. E.g., at $\mathcal{O}(\alpha^{0})$
$G_{C}^{(0)}(z)=i\int_{0}^{\infty}~{}\text{d}u~{}\chi_{A}(-u)~{}e^{-iuz}~{}=G_{A}(z)~{}.$ (34)
@@
Therefore, all orders of correction can be obtained in terms $G_{C}^{(0)}(z)=G_{A}(z)$ (and its various order derivatives with respect to $z$), which in tern can be obtained from the characteristics function of the distribution $\rho_{A}(\lambda)$. Note also that due to the specific scaling law for the $R$-transformation, the $n$-order correction to $G_{C}^{(0)}(z)$ is proportional to $\alpha^{2n}$. The approximate DOS of the operator $C$ can now be obtained from $G_{C}(z)$ using the Stieltjes inversion formula (7),
$\rho_{C}(\lambda)=\frac{1}{\pi}\lim_{\epsilon\rightarrow 0}(\text{Im}~{}G_{C}(\lambda-i\epsilon))~{}.$ (37)

### v2

@@ Abstract
Utilizing the framework of free probability, we analyze the spectral and operator statistics of the Rosenzweig–Porter random matrix ensembles, which exhibit a rich phase structure encompassing ergodic, fractal, and localized regimes. Leveraging subordination formulae, we develop a perturbative scheme that yields semi-analytic expressions for the density of states up to second order in system size, in good agreement with numerical results. We compute higher-point correlation functions in the ergodic regime using both numerical and suitable analytic approximations. Our analysis of operator statistics for various spin operators across these regimes reveals close agreement with free probability predictions in the ergodic phase, in contrast to persistent deviations observed in the fractal and localized phases, even at late times. Notably, the fractal phase exhibits partial freeness while retaining memory of the initial spectrum, highlighting the importance of non-localized eigenstates and associated with the late-time dynamics of cumulative out-of-time-ordered-correlators (OTOCs). Employing distance measures and statistical tools such as the $\chi^{2}$ statistic, Kullback–Leibler divergence, and Kolmogorov–Smirnov hypothesis testing, we define a characteristic time scale – the free time – that marks the onset of the validity of free probability predictions for operator spectral statistics in the ergodic phase. Remarkably, our findings demonstrate consistency across these different approaches.
@@ 1 Introduction
Although similar techniques have been previously applied to analyze the DOS in various models under constrained settings, our method has the practical advantage that it directly uses the free probability results (which otherwise one usually needs to establish first, e.g., using the replica technique and so on) and is also valid for a general class of Hamiltonians where one of the two terms of the Hamiltonian is Gaussian random matrix. As one of the main results of the paper, we obtain the first two orders of corrections to the Gaussian distribution of the DOS of the RP model when the system is in the fractal phase. By numerical analysis, we show that our approximate analytical expression for the DOS is in very good agreement with it.
We further study the statistics of the sum of two deterministic spin operators, one of which is time-evolved by the RP model Hamiltonian. At the initial stage, the eigenvalue distribution of the total operator adheres to the free binomial distribution. However, as time progresses, the eigenvalue distribution exhibits significant deviations from this initial distribution. Specifically, in the ergodic regime, the eigenvalue statistics of the operator sum show a transition from the multinomial distribution to the one predicted by the free probability distribution when the component operators are free. The multinomial distribution (discussed later) is the general case of the binomial distribution, which holds only for spin-half operators. Conversely, in the fractal and localized regimes, the eigenvalue distributions diverge substantially from the predictions obtained from free probability theory. This observation lends credence to the broader assertion that chaos leads to the asymptotic freeness between two operators at late times. The behavior of operator statistics is consistent with the vanishing contribution of the cumulative higher-point OTOCs at late times.
@@ 2 Elements of Free Probability theory
Having discussed the condition that two mutually free random variables, say $a$ and $b$, should satisfy, [^3] Throughout this paper, we assume that these two variables are self-adjoint and have well-defined eigenvalue density. we now describe the procedure for obtaining the free probability prediction for the density of eigenvalues of the sum ($a+b$) of these two operators, which will be used in Sec. 6 to compare with numerically obtained distributions. The eigenvalue distribution of $a+b$ is called the free convolution of the eigenvalue distributions $\rho_{a}(\lambda)$ and $\rho_{b}(\lambda)$, and is usually denoted as $\rho_{a+b}(\lambda)=\rho_{a}(\lambda)\boxplus\rho_{b}(\lambda)$ nica2006lectures .
The first ingredient we need to define is the resolvent operator associated with the operators $a$ and $b$, which is the Cauchy transformation of the DOS of the respective operator, e.g.,
@@ 2.1 Summary of Known Results on Free Additive Convolution
For operators with spin larger than unity, it becomes increasingly challenging to obtain analytic formulae for the free additive convolution of the sum of operators, due to the fact that one needs to solve order-$s$ algebraic equations. Fortunately, one can obtain such distributions numerically using subordination function methods; see, for instance speicher2019lecturenotesfreeprobability . Such methods were applied in Camargo:2025zxr for the derivation of the free additive convolution of spin-$3/2$ and spin-$2$ operators. Figure 1 shows the respective distribution for spin-$1/2$, spin-$1$, and spin-$3/2$. We review the subordination function method in Appendix B.
@@ 2.2 Asymptotic Freeness Produced by Chaotic Dynamics
Given two operators $P$ and $Q$ that are not initially free, the time evolution of $Q$ under a chaotic Hamiltonian is expected to produce a Heisenberg operator $Q(t)=e^{iHt}\,Q\,e^{-iHt}$ that becomes free from $P$ Jindal:2024zcg ; Chen:2024zfj . This emergence of freeness defines a novel notion of quantum chaos, which can be incorporated into the quantum ergodic hierarchy Camargo:2025zxr .
Freeness arises only in infinite-dimensional Hilbert spaces. When operators $P$ and $Q$ are represented as $N\times N$ matrices, freeness can emerge only in the limit $N\to\infty$, a regime associated with asymptotic freeness. For finite $N$, cumulants never vanish completely and retain residual values at late times. However, under chaotic dynamics – i.e., when the Hamiltonian exhibits features akin to random matrix theory – the magnitude of these residual cumulants decreases with increasing $N$, suggesting their vanishing in the $N\to\infty$ limit. In this sense, the finite-$N$ system “knows” about its large-$N$ counterpart.
This connection also manifests in operator statistics. Although true freeness between $P$ and the time-evolved operator $Q(t)$ requires infinite $N$, the spectral statistics of $P+Q(t)$ at finite $N$ often agree with free probability predictions. This agreement serves as a practical signature – or “smoking gun” – of asymptotic freeness.
2.3 Relation with $2n$-point functions
[…]
When one of these operators is a time-evolved Heisenberg operator (as is the case we consider below in section 6), the moments $\langle(P+Q(t)^{k}\rangle$ have expansions in terms of the OTOCs of different orders. [^5] We thank Koji Hashimoto for discussions on this point. To see this explicitly, consider two spin-$1/2$ operators and the following second-order moment:
$\langle(P+Q(t))^{2}\rangle=\langle P^{2}\rangle+\langle Q(t)^{2}\rangle+2\langle PQ(t)\rangle=2+2\langle PQ(t)\rangle\,,$ (16)
where the normalization $\langle P^{2k}\rangle=\langle Q^{2k}\rangle=1$ has been applied throughout, and the operators are chosen such that their one-point function vanishes: $\langle P\rangle=\langle Q(t)\rangle=0$. In other words, we choose the traceless operators or the traceless part of the operators by appropriately subtracting the traces. For the random matrix ensembles, the average is taken over the ensembles of the Hamiltonians.
Similarly, for the fourth-order moment, we have:
$\langle(P+Q(t))^{4}\rangle=6+2\langle PQ(t)PQ(t)\rangle\,,$ (17)
where by definition $\langle P^{4}\rangle=\langle Q^{4}\rangle=1$. In this expression, the second term can be identified as the lowest-order OTOC. Consequently, Eqs. (16) and (17) reveal that the moments of $P+Q(t)$ encode information of both the two- and four-point OTOCs.
These relationships can be extended to higher-order moments. In chaotic systems, at sufficiently late times, the time evolution is expected to render deterministic operators asymptotically free. Consequently, for operators with vanishing one-point functions, the mixed moments vanish. As a result, the even-order moments simplify under the conditions $\langle P^{2n}\rangle=\langle Q^{2n}\rangle=1$ and $\langle P^{2n+1}\rangle=\langle Q^{2n+1}\rangle=0$. Therefore, one obtains
$\langle(P+Q(t))^{2n}\rangle\simeq\binom{2n}{n}\,,~{}~{}~{}~{}~{}t\rightarrow\infty\,,$ (18)
while all odd moments vanish.
As expected, these moments correspond directly to those of the free probability prediction for the eigenvalue distribution of the operator $P+Q(t)$ at late times. For example, it is easy to see that the moments of the arcsine distribution in (12) are given by the same binomial coefficients:
$\int_{-2}^{2}\mathrm{d}\lambda~{}\frac{\lambda^{2n}}{\pi\sqrt{4-\lambda^{2}}}=\binom{2n}{n}\,.$ (19)
Thus, for the sum of two spin-$1/2$ operators, the eigenvalue distribution converges to the arcsine law predicted by the free probability. Extensions to higher spins follow naturally.
We emphasize that this connection is not a mere coincidence; rather is a direct consequence of the fact that when the time evolution generates free variables, the moments of an operator of the form $P+Q(t)$ should approach at late times those of the distribution predicted by the free additive convolution of the eigenvalue distributions of those of $P$ and $Q$.
@@ 4.1 Cauchy and R-transforms
From the functional inverse of the resolvent of an operator $\mathcal{O}$ (i.e., the Cauchy transformation of the density of eigenvalues of the operator $\mathcal{O}$), we define the $R$-transformation
$G_{\mathcal{O}}^{-1}(z)\equiv\mathcal{B}_{\mathcal{O}}(z)=\frac{1}{z}+R_{\mathcal{O}}(z)~{}.$ (24)
We note that the expression for the resolvent of the operator $A$ in terms of its DOS, $\rho_{\mathcal{O}}(\lambda)$ as
$G_{\mathcal{O}}(z)=\int\frac{\rho_{\mathcal{O}}(\lambda)}{z-\lambda}~{}\text{d}\lambda~{}.$ (25)
Here, the integral is over the support of the distribution $\rho_{\mathcal{O}}(\lambda)$. Note that $G_{\mathcal{O}}(z)$ is essentially the Cauchy transformation of $\rho_{\mathcal{O}}(\lambda)$.
When the operator $\mathcal{O}$ is the Hamiltonian of some quantum mechanical system belonging to some random matrix ensemble, one can relate the resolvent with the time evolution operator as follows. Consider the trace of the ensemble-averaged time evolution operator,
@@
$\rho_{H}(\lambda)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\text{d}t~{}e^{-it\lambda}~{}\mathcal{U}_{H}(t)~{}.$ (27)
@@ 4.2 Addition of two free operators: subordination formulae for the resolvent
Now, for addition of two such operators $P$ and $Q$, which are free, from the additivity of the $R$ transformation, we have, for the operator $C=P+Q$, the following,
@@
$\mathcal{B}_{P}\big{(}G_{C}(z)\big{)}=z-R_{Q}\big{(}G_{C}(z)\big{)}\,.$ (30)
Finally, applying $G_{P}$ on both sides, we get a self-consistency equation for $G_{C}(z)$,
$G_{C}(z)=G_{P}\big{(}z-R_{Q}\big{(}G_{C}(z)\big{)}\big{)}\,.$ (31)
One can perform an entirely analogous procedure for the operator $Q$ and obtain
$G_{C}(z)=G_{Q}\big{(}z-R_{P}\big{(}G_{C}(z)\big{)}\big{)}~{}.$ (32)
In the following, we use these formulae to obtain an expression for the DOS for the RP model. The choice between the two formulae for obtaining the density of eigenvalues of $C$ – which can be extracted from $G_{C}(z)$ using the Stieltjes inversion formula – depends on whether the resolvent and the $R$-transformation are more straightforward for the operators $P$ or $Q$.
Before proceeding further, we note that a self-consistency equation analogous to the one in (31) for the resolvent of the RP model was obtained in Venturelli:2022hka using the replica trick. The authors also discussed the connection of their formula with the one obtained from the law of addition in free probability. The discussion on the approximate evaluation of the DOS of the RP model below follows closely the discussion of Ref. Venturelli:2022hka . However, we have kept the discussion below general enough so that it applies not only for the RP model, but addition of two non-commutative operators, one of which has the Wigner semicircle as the DOS.
@@ 4.3 A perturbative scheme for the resolvent
For most of the choices of the operators $P$ and $Q$, the expressions for the corresponding resolvents are sufficiently complicated so that it is difficult to solve either of the equations in eqs. (31) or (32) for $G_{C}(z)$. Therefore, we now proceed as follows to obtain a perturbative expression for the DOS of an operator of the form $C=P+\alpha Q$, where $P$ and $Q$ are mutually free and the expressions for their DOS are assumed to be known. Here $\alpha=N^{-\gamma/2}$ is a parameter which we assume to be small so that one can obtain a perturbation series. Using the scaling transformation rule of the $R$-transformation, [^8] The $R$-transformation of the density of eigenvalues of an operator $P$ follows the scaling law: $R_{\alpha P}(z)=\alpha R_{P}(\alpha z)$, where $\alpha$ is a constant. and the definition (in Eq. (25)), we have, from Eq. (31),
$G_{C}(z)$ $=\int\text{d}\lambda~{}\frac{\rho_{P}(\lambda)}{z-\alpha R_{Q}\big{(}\alpha G_{C}(z)\big{)}-\lambda}$
$=i\int_{0}^{\infty}~{}\text{d}u~{}\chi_{P}(-u)\exp[-iu\Big{(}z-\alpha R_{Q}\big{(}\alpha G_{C}(z)\big{)}\Big{)}\Big{]}~{}.$ (33)
Here we have defined the characteristics function $\chi_{P}(u)$ of the distribution $\rho_{P}(\lambda)$ by a Fourier transformation
$\chi_{P}(u)=\int\text{d}\lambda~{}\rho_{P}(\lambda)~{}e^{-i\lambda u}~{}.$ (34)
Comparing this expression with the inverse of Eq. (27), we see that the characteristic function is nothing but the ensemble-averaged evolution operator generated by $P$, i.e., $\chi_{P}(-t)=\mathcal{U}_{P}(t)$.
@@
$G_{C}(z)=i\int_{0}^{\infty}~{}\text{d}u~{}\chi_{P}(-u)~{}e^{-iuz}\sum_{n}\frac{(iu\alpha)^{n}}{n!}\big{(}R_{Q}\big{(}\alpha G_{C}(z)\big{)}\big{)}^{n}\,.$ (35)
Writing in terms of the free cumulants ($\kappa_{j}(Q)$) of the operator $Q$, this has the form
$G_{C}(z)=\sum_{n}\frac{i^{n+1}\alpha^{n}}{n!}\Big{(}\sum_{j=1}^{\infty}\kappa_{j}(Q)\big{(}\alpha G_{C}(z)\big{)}^{j}\Big{)}^{n}\int_{0}^{\infty}~{}\text{d}u~{}\chi_{P}(-u)~{}u^{n}~{}e^{-iuz}\,.$ (36)
So far, this expression for $G_{c}(z)$ is exact. For a general operator $Q$, its $R$-transformation can be an arbitrary function of $z$, so that it is difficult to obtain a perturbation series with $\alpha$ as a small parameter. However, one particular case where the above expression becomes easily tractable is when the operator $B$ has Wigner semicircle as the DOS (which is the case also for the RP model considered in this paper), i.e., $\rho_{Q}(\lambda)=\frac{1}{2\pi}\sqrt{4-\lambda^{2}}$. In that case [^9] The Wigner semicircle is the analogue of the Gaussian distribution for non-commutative probability, i.e., this distribution has only the first free cumulant non-zero. Since $R(z)$ is the generator of the free cumulants, this is equal to $R(z)=z$ for the semicircle distribution. $R(z)=z$, and the expression in Eq. (35) (or (36)) simplifies to
$G_{C}(z)$ $=\sum_{n}\frac{i^{n+1}}{n!}\alpha^{2n}\big{(}G_{C}(z)\big{)}^{n}\int_{0}^{\infty}\text{d}u~{}u^{n}\chi_{P}(-u)e^{-iuz}\,,$
$=\sum_{n}\frac{(-1)^{n}}{n!}\alpha^{2n}\big{(}G_{C}(z)\big{)}^{n}\partial_{z}^{n}G_{P}(z)\,,$ (37)
where to obtain the last expression, we have used an analogous relation to that of the one in (28), valid for the operator $P$.
Evaluating the terms of the above sum recursively, one can obtain the resolvent $G_{C}(z)$ with corrections of different orders of $\alpha$. Thus, at $\mathcal{O}(\alpha^{0})$, we have
$G_{C}^{(0)}(z)=i\int_{0}^{\infty}~{}\text{d}u~{}\chi_{P}(-u)~{}e^{-iuz}~{}=G_{P}(z)\,.$ (38)
@@
Therefore, all orders of correction can be obtained in terms $G_{C}^{(0)}(z)=G_{P}(z)$ (and its various order derivatives with respect to $z$), which in tern can be obtained from the characteristics function of the distribution $\rho_{P}(\lambda)$. Note also that due to the specific scaling law for the $R$-transformation, the $n$-order correction to $G_{C}^{(0)}(z)$ is proportional to $\alpha^{2n}$. The approximate DOS of the operator $C$ can now be obtained from $G_{C}(z)$ using the Stieltjes inversion formula (7),
$\rho_{C}(\lambda)=\frac{1}{\pi}\lim_{\epsilon\rightarrow 0}(\text{Im}~{}G_{C}(\lambda-i\epsilon))\,.$ (41)